When R=10, V=15, when R=20,V=20 What is vif R is infinite? M R * en s
Answer #1

0.00 Date: / V-V + Ovo e REV R(V-V-) = VR, - when R = lor, v=15w 10(15-V1) = 15 R- 15-V7 = 115 R — when R=200 v=20v 20(20-V-)Dat I from O R(V-V.) = VRT RV-RV = VRT V(R-RI) = RVT v R VE R- R7 = RVT Put R=0 v7 = 10 R=10 v=lov I